OpenStudy (anonymous):

A game plan to use a variation of the Tarot deck for character creation. The idea is to have a card randomly selected for each of the 6 character traits, and assigning a score and special trait based on the card selected. If the game only uses 22 cards Major Arcana, how many different characters can be created?

OpenStudy (anonymous):

do you have ur combination and permutation equations?

OpenStudy (anonymous):

This seems to be like a very difficult problem to me.

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

I have trying to figure out how to set that up DemolisionWolf

OpenStudy (anonymous):

its been a while since i've done these so give me a min. and if anyone else remembers, fell free to post ^_^

OpenStudy (anonymous):

Alright I will read over the question again to see if I understand it more.

OpenStudy (anonymous):

are the cards placed back in the deck after the character has been made, or are the cards left out after each character is made?

OpenStudy (anonymous):

I believe the cards are left out

OpenStudy (anonymous):

On my assignment it says the answer is really big

OpenStudy (anonymous):

if the cards are left out then at the end of character 1, we would have 22-6 cards remaing = 16 at the end of character 2, we would have 16-6=10 cards remaining at the end of character 3, we would have 10-6=4 cards remaining. so taking that route would leave us 3 characters created. If the cards are put back in then..

OpenStudy (anonymous):

oh does it? ok, then I think I know what to do, we have to use an exponenet-

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

since you haven't learned about combinations and permutations, then this is the equation you want to use: \[NumberOfCards^{NumberOfTraits} = TotalPossibleCharacters\]

OpenStudy (anonymous):

I have learned about permutations just not good at them

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I like the formula you came up with

OpenStudy (anonymous):

haha, so that is the permutation formula, it figures out all the possible ways that 22 cards can be ordered into groups of 6. so did you have 22^6 as your equation?

OpenStudy (anonymous):

No I am still trying to come up with my equation

OpenStudy (anonymous):

How did you come up with that equation

OpenStudy (anonymous):

are u there still

OpenStudy (anonymous):

so the equation is nothing special, it is just a way to say how many possible ways there are to organize 22 cards in groups of 6 cards. \[n^r\] where r is the number of 'slots' that are possible and n is the number of total options available. for example, if we had a combination lock, and it had one dial on it and the dial goes from 0 to 9, then, the number of 'slots' would be 1, becuause there is only one time we can have a number. and the number of 'options' would be 10, becuase we go from 0 to 9. so the equation would look like: 10^1 = 10, which is total number of possible combinations to have the lock set to. which makes sense to you and me. imagine a combination lock that only has an option to be 0, 1, 2, 3, 4, 5, 6, 7, 8, and 9. then there are 10 total ways possible. I hope this helps I have to go now

OpenStudy (anonymous):

10 is the answer

OpenStudy (anonymous):

the answer to your question is 10? hmm...

OpenStudy (anonymous):

no I was asking you what you got in the example

OpenStudy (anonymous):

oh ok. ya, 10 is what it would be in the example.

OpenStudy (anonymous):

How would I get the answer in my problem?

OpenStudy (anonymous):

\[22^6\]

OpenStudy (anonymous):

Yeah that what I already have the answer for that problem